# Forcing Planets to Evolve: The Relationship Between Uranus and Neptune at Late Stages of Dynamical Evolution

### v1

@@ 4.1 Reproducing Epochs of Migration and Eccentricity Damping
$\frac{de_{P}}{dt}=-\frac{\pi}{\sqrt{2}}\frac{G}{\Omega}\frac{M_{P}}{M_{*}}\ln(1+\Lambda^{2})\frac{\Sigma}{H}\frac{1}{e_{l}^{2}+e_{P}^{2}}$ (13)
$t_{\rm damp}=\frac{e_{P}}{de_{P}/dt}$ (14)
@@
Plugging in the appropriate disk and Neptune parameters from Tsiganis et al. (2005) simulation and scaling the mass volume density of the disk due to the dispersal after millions of years, we calculate one million years and eight million years for Neptune’s early and late damping phases, respectively. These values match the damping timescales from the simulation (see Figure 5). The physical process of dynamical friction is responsible for damping Neptune’s eccentricity in the simulation, and it turns out to be well-represented by an exponential function with an order of magnitude estimated timescale.
We confirm the appropriate use of an exponential function representing the dynamical friction process due to interactions between the planets and massive planetesimals. We use the best-fit values to initialize simulations of the giant planets with migration and eccentricity damping parameters. In Figure 6 we show a simulation where we successfully reproduce early and late epochs of outward migration (top panel) and eccentricity damping for Uranus (middle panel) and Neptune (bottom panel). With this exercise, we show an appropriate use for aritficially evolving planets’ orbits, motivated by a physical process. While for this case, our goal was to reproduce the scenario is a specific simulation, for the case of outer solar system models, the goal is to evolve the giant planets’ orbit such that they match the current observed state of the Solar System. This involves changing the parameters in the functional form describing the evolution. Specifically, the parameters to change are the scale of the change and the timescale of the change (see Equation 7).
@@
Figure 6: Two example simulations (orange) showing the evolution of the semimajor axes (top panel) and eccentricities of Uranus (middle panel) and Neptune (bottom panel) representing an early and late damping phase (separated by the shaded region). We artificially evolve the two outer planets’ orbits using the best-fit parameters from Figure 5. The evolution from Figure 5 is shown for reference in blue. The dashed curves show the best fit exponential evolution for each early and late migration/damping phase. We note that because this is a chaotic problem, slightly different initial conditions produce vastly different results. Simply randomizing the initial phase angles from 0 to $2\pi$ for an ensemble of 40 simulations produced $\approx$ 20 simulations with completely different final outcomes.
@@ 4.2 Effect of Damping Neptune on Uranus
An interesting outcome of our code tests is that of isolating the impact of exclusively damping Neptune’s eccentricity as if it had been damped by dynamical friction. With this method, we study how the secular coupling between Uranus and Neptune at early and late stages affects the evolution of Uranus. This effect is similar to the redistribution of tidal damping by secular coupling between planets (Greenberg et al., 2013). By modeling Neptune’s evolution artificially and without a massive planetesimal disk, we can study the interactions between Uranus and Neptune in isolation.
[…]
@@
Figure 7: Fifty million year simulation initialized with the same parameters as those in the late stage evolution from Figure 6 and only showing Uranus (top row) and Neptune (bottom row). Neptune’s eccentricity is damped artificially with the functional form shown in dark maroon (bottom middle panel). Uranus’ is not artificially damped in this simulation, rather, it is indirectly damped (top middle panel). The dashed line shows the same damping timescale for Uranus as in Figure 6 to show that the indirect damping of Uranus is comparable to that calculated from Tsiganis et al. (2005).
@@
We run additional simulations similar to those shown in Figure 7, but we vary the initial eccentricity of Neptune and Uranus to investigate how changing the $\Delta e$ will affect the indirect damping on Uranus. Increasing the initial eccentricity of both Uranus and Neptune to a higher value and damping Neptune with the same timescale of 6 million years, results in Uranus’ eccentricity damping to a higher degree (not shown). This effect is interesting to follow up on when studying the late stages of planetary evolution in our solar system. The secular eccentricity timescales of Uranus and Neptune are 0.5 Myr and 2 Myr, therefore the secular coupling between Uranus and Neptune, that would cause Neptune to damp Uranus’ eccentricity indirectly, is stronger at late times when the damping timescale is $\ga 6$ Myr.
@@
While the giant planet instability model has reproduced a vast amount of currently observed structure within the giant planets and Trans-Neptunian Objects, the evolution in between an instability and the current TNO configuration remains unknown. This includes the possibility that Uranus and Neptune were launched onto orbits with eccentricities as high as 0.4 and damped at various magnitudes and rates depending on the volume density of the planetesimal disk. With this in mind, we run a last suite of simulations similar to those described in Figure 7 with Uranus’ eccentricity initialized at values higher than Neptune’s. One such example is shown in Figure 8. In this simulation, Uranus’ and Neptune’s eccentricity are initialized at 0.1 and 0.065, respectively, representing a scenario just before a damping phase. Both planets are artificially damped with the same eccentricity-damping timescale, motivated by instability simulations that show the number density near Uranus and Neptune is comparable after a few million years post-instability (not shown). We find that Neptune’s effect on Uranus is significant enough to damp Uranus’ eccentricity by a larger amount than what we prescribed (top middle panel). The secular coupling between Uranus and Neptune at this late stage of evolution is important to understand in order to constrain how dynamical friction sets the final eccentricities of the ice giants.
@@
Figure 8: Fifty million year simulation initialized with the same parameters as those in the late stage evolution from Figure 6, except Uranus (top row) and Neptune (bottom row) start at eccentricities 0.1 and 0.065, respectively. Both planets are artificially damped with the same timescale of 5 Myr. Uranus’ eccentricity damps further than the amount we prescribe (top middle panel) due to the secular effects with Neptune.
4.3 Discussion of Future Applications
In this paper we focus on Uranus’ and Neptune’s artificial orbital evolution in the context of reproducing migration/eccentricity damping phases in a well-known planetary instability simulation and isolating the impact of one planet’s eccentricity damping in the absence of a planetesimal disk. Artificial evolution is also useful for outer solar system studies that probe how different giant planet dynamical histories affect the orbital configuration of Trans Neptunian Objects (TNOs). A variety of proposed migration histories within the outer solar system, modeled through numerical simulations, produce different planet and TNO structure (for reveiws see Nesvorný, 2018; Gladman & Volk, 2021). In fact, a similar migration history, with minuscule changes in initial conditions, could also produce vastly distinct planet trajectories as is the case for planetary instability numerical models as discussed in Section 4, where chaos is the underlying driver Volk & Malhotra (2019).
A realistic portrayal of a planetary system in an N-body simulation would require all solar system planets and billions of massive planetesimals. The chaotic nature of this problem is further accentuated by the additional gravitational interactions between each body, making exploration of a large parameter space of initial conditions computationally infeasible. The ability to artificially evolve planets’ orbits is therefore beneficial for investigating how TNOs in mean motion resonance (MMR) with Neptune evolve as a result of various migration histories.
5 Summary
@@
The modifications to REBOUND and Mercury6.2 we present in this paper are essential for running suites of N-body simulations that can probe a large parameter space of orbital configurations for the purpose of constraining the evolution of the outer solar system. We demonstrate how to evolve particles’ orbital evolutions with arbitrary functions using our modifications. We show that with a one-planet simulation, all orbital elements can be evolved independently from each other. Introducing a second planet to the system, adds new gravitational interactions between the planets, which affects the amount of control one has over the artificial evolution of each orbital element. Secular perturbations between the planets produce oscillations in eccentricity and inclination. It is important to consider how the secular frequencies and amplitudes of the planets compare to the desired scale of change and timescale of change to successfully evolve these elements accordingly.
We demonstrate two appropriate uses of this code, emphasizing the need for physically motivated functional forms for a given orbital evolution. In our case, we used the planetary instability simulation in Tsiganis et al. (2005) to validate the use of our code and artificially evolve the semimajor axis and eccentricity of Uranus and Neptune. We choose an exponential function to evolve these parameters and confirm migration/damping timescales with dynamical friction derivations appropriate for the planetesimal disk in their simulation. We are able to reproduce two stages of migration and eccentricity damping, which we call the early and late stages. This result is important for studying various migration scenarios post-instability and how that affects the orbital distribution of trans-neptunian objects. Lastly, we find that artificially damping Neptune’s eccentricity indirectly damps Uranus’ eccentricity due to their secular coupling. At late stages of evolution, coupling between the latter two planets is important to consider since the planetesimal disk will have dissipated significantly.
Acknowledgements
@@
We gratefully acknowledge funding from NASA Emerging Worlds grant 80NSSC21K0376. AHR thanks the LSST-DA Data Science Fellowship Program, which is funded by LSST-DA, the Brinson Foundation, and the Moore Foundation; her participation in the program has benefited this work. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 1842400. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. KV acknowledges additional funding from NASA grants 80NSSC19K0785 and 80NSSC23K0886. RP acknowledges additional funding from NASA Solar System Observations grant 80NSSC21K0289. We acknowledge use of the lux supercomputer at UC Santa Cruz, funded by NSF MRI grant AST 1828315.
Data Availability
REBOUND code, and example scripts to create the figures from this paper are available on my github: https://github.com/ahermosillo. Simulation data is available upon request.
Appendix A Modifications to the equations of motion

### v2

@@ 4.1.2 Calculating Eccentricity Damping Timescale due to Dynamical Friction
$\frac{de_{P}}{dt}=-\frac{4\pi}{3\sqrt{3}}\left(\frac{\Sigma a_{P}^{2}M_{P}}{M_{*}^{2}}\right)\left(\frac{e_{P}}{e_{l}^{4}}\right)\tilde{\Omega}_{P}\ln(1+\Lambda^{2})$ (9)
where $\Sigma=(m_{tot}/(2\pi r_{out}(r_{out}-r_{in}))(r_{out}/a_{P})$, $\rho=\Sigma/(2H)$, and $H=a_{P}\sin{i_{l}}\approx a_{P}e_{l}/\sqrt{2}$. The characteristic order-of-magnitude damping timescale due to dynamical friction is then
$t_{\rm damp}=\frac{e_{P}}{de_{P}/dt}.$ (10)
where $de_{P}/dt$ is evaluated using Equation (9).
@@
For Neptune we plug in $M_{P}=M_{\rm N}$, $a_{P}=25.5$ au for the early ($e_{P}=0.06$) and late ($e_{P}=0.055$) damping phases. For Uranus, we plug in $M_{P}=M_{\rm U}$, $a_{P}=19.3$ au for the early ($e_{P}=0.06$) and late ($e_{P}=0.054$) phases. For both cases, we plug in $M_{*}=M_{\odot}$, $e_{l}=e_{P}$, $m_{tot}=35M_{\oplus}$, $r_{out}=30$ au, and $r_{in}=18$ au. Then, we scale the mass density, $\rho=\Sigma/(2H)$, to account for the dispersal after millions of years by dividing $\rho$ by 50 and 500 for the early and late damping phases, respectively. We calculate $\approx 0.8$ Myr and $\approx 6.1$ Myr for Neptune’s early and late damping phases, respectively. For Uranus, we calculate $\approx 0.7$ Myr and $\approx 5.5$ Myr. These values match the damping timescales from the simulation within a factor of 2 (see Figure 5) with the exception of the late damping phase of Uranus, possibly suggesting that another process is in effect here (see Section 4.2). To estimate the normalizing factor for the mass density, which accounts for the disk dissipation over time, we tracked how the number density near Neptune evolved in an example dynamical upheaval simulation. In comparing the calculated damping timescale with the best-fit simulation timescale, we find that the physical process of dynamical friction is responsible for damping Neptune’s eccentricity in the simulation, and it turns out to be well-represented by an exponential function with an order of magnitude estimated timescale.
We don’t reproduce the final asymptotic eccentricities with this formulation. Without the mass density of the disk as a function of time, we are unable to integrate Equation (9). Uranus’ final eccentricity is higher than Neptune’s, possibly because the mass density around it dissipated faster, but the simulation data needed to test this hypothesis is unavailable. The planets’ eccentricities eventually asymptote because the overall disk’s mass density decreases significantly over time.
@@
Figure 6: Two example simulations (orange) showing the evolution of the semimajor axes (top panel) and eccentricities of Uranus (middle panel) and Neptune (bottom panel) representing an early and late damping phase (separated by the shaded region). Initial conditions and imposed evolution are found in Table 1. We artificially evolve the two outer planets’ orbits using the best-fit parameters discussed in Section 4.1.1. The evolution from Figure 5 is shown for reference in blue. The dashed curves show the best fit exponential evolution for each early and late smooth migration and eccentricity damping phases.
4.1.3 Simulating the Early and Late Migration and Eccentricity Damping Phases
Having confirmed that the use of an exponential decay function to represent the dynamical friction process is appropriate, we use the best-fit values described in Section 4.1.1 to artificially migrate and damp Uranus and Neptune. In Figure 6 we show a simulation where we successfully reproduce early and late epochs of outward migration (top panel) and eccentricity damping for Uranus (middle panel) and Neptune (bottom panel). The initial planet parameters and imposed evolution are shown in Table 1. We note that because this is a chaotic problem, slightly different initial conditions produce vastly different results. For example, a minuscule change in argument of pericenter may enact a different gravitational interaction between the planets that causes one planet to be ejected from the system. We ran 40 instances of these simulations with the same parameters as described in Figure 6, except the phase angles ($\omega$, $\Omega$, and $f$) were initialized randomly from a range of 0 to $2\pi$. We saw two different outcomes: in 20 simulations, the planets behaved similarly to those in Figure 6, and in the remaining simulations, the planets underwent a chaotic evolution where Uranus and/or Neptune reached high semi-major axis/eccentricities or were ejected.
With this exercise, we show an appropriate use for aritficially evolving planets’ orbits, motivated by a physical process. While for this case, our goal was to reproduce the scenario in a specific simulation, for the case of outer solar system models, the goal is to evolve the giant planets’ orbits such that they match the current observed state of the Solar System. To do this, one would use a function and parameters that represents a realistic temporal evolution (see Equation 7).
@@ 4.2 Effect of Damping Neptune on Uranus
An interesting outcome of our code tests is that of isolating the impact of exclusively damping Neptune’s eccentricity as if it had been damped by dynamical friction. With this method, we study how the secular coupling between Uranus and Neptune at early and late stages affects the evolution of Uranus, when only artificially evolving Neptune. This effect is similar to the redistribution of tidal damping by secular coupling between planets (Greenberg et al., 2013). By modeling Neptune’s evolution artificially and without a massive planetesimal disk, we can study the interactions between Uranus and Neptune in isolation.
We find that as Neptune’s eccentricity damps, Neptune exerts a force on Uranus, such that Uranus’ eccentricity damps as well. The magnitude of the effect depends on a few parameters. If Neptune’s damping timescale is short ($\lesssim$ 1 Myr) or the change in eccentricity is small ($\lesssim$ 0.01), then the effect on Uranus is negligible. We run simulations with a variety of initial eccentricities for Uranus and Neptune and damping timescales to understand this effect. All simulations contain the 4 giant planets with initial parameters and imposed evolution described in Table 1. We show three examples in the following text with parameters described in Table 1. Similar to the simulations described in Figure 6, we ran 40 simulation instances of each example, where we randomly chose the phase angles ($\omega$, $\Omega$, $f$) between 0 and $2\pi$.
Figure 7 shows a simulation that is initialized in the same way as the late-phase simulation from Figure 6, but only Neptune’s eccentricity and semimajor axis are artificially evolved (bottom panels; maroon lines). We keep Uranus’ semi-major axis fixed since it only migrated a fraction of an au, and its migration does not affect the eccentricity evolution here. Neptune is expected to migrate a larger distance, and so we represent that in these simulations. Through experiments, we found that this coupling behavior is independent of the migration of both planets. We find that Uranus’ eccentricity is indirectly damped in this simulation through interactions with Neptune (top right panel). Uranus damps to a similar extent to the evolutionary behavior seen in the simulations from Tsiganis et al. (2005). For reference, we plot the same exponential damping curve from Figure 6 on top ($\tau_{e}=20Myr$, $e_{0}=0.054$, $\Delta e=-0.024$). In the 40 simulations we ran, we found two outcomes: $12$ simulations behaved similarly to Figure 7 and the remaining simulations showed chaotic planet evolution.
@@
Figure 7: Fifty million year simulation (orange dots) initialized with the same parameters as those in the late stage evolution from Figure 6 (also, see Table 1) and only showing Uranus (top row) and Neptune (bottom row). Neptune migrates and its eccentricity is damped with the functional form shown in maroon (bottom panels). Uranus’ is not artificially damped in this simulation, rather, it is indirectly damped (top right panel). The dashed line shows the same damping timescale, 20 Myr, for Uranus as in Figure 6 to show that the indirect damping of Uranus is comparable to that from Tsiganis et al. (2005).
@@
We ran additional simulations similar to those shown in Figure 7, but we vary the initial eccentricity of Neptune and Uranus to investigate how a different $\Delta e$ and $\tau_{e}$ will affect the indirect damping on Uranus. We show one example in Figure 8. If the initial eccentricity of both Uranus and Neptune are larger ($0.06$ and $0.065$, respectively) and Neptune’s damping timescale is shorter ($\tau_{e}=5.1$ Myr), Uranus’ eccentricity damps a larger amount and faster than in the example from Figure 7. We saw two outcomes in the 40 simulation instances: $12$ simulations behaved similarly to Figure 8 and the remaining simulations showed chaotic planet evolution. This effect is interesting to follow up on when studying the late stages of planetary evolution in our solar system where the exact details of the planets’ eccentricities and planetesimal remain uncertain. In a simulation representing the early phase described in Figure 6, we only exponentially damped Neptune’s eccentricity from 0.06 to 0.02, where $\tau_{e}=1.5$ Myr and saw no effect on Uranus (plot not shown). The damping timescale is too short at this early phase (i.e., it is comparable to and shorter than the secular frequency of Uranus and Neptune, respectively), so Neptune’s eccentricity damping does not indirectly damp Uranus. The secular eccentricity timescales of Uranus and Neptune are 0.5 Myr and 2 Myr, respectively (see, e.g. Murray & Dermott, 1999); therefore, the secular coupling between Uranus and Neptune, that would cause Neptune to damp Uranus’ eccentricity indirectly, is stronger at late times when the damping timescale is $\gtrsim 6$ Myr.
Figure 8: Fifty million year simulation similar to that in Figure 7, except Uranus and Neptune have higher initial eccentricities of 0.06 and 0.065 and Neptune’s eccentricity is damped with a timescale of 5.1 Myr. For reference two dashed lines are plotted on top of Uranus’ eccentricity (top right) with $\tau_{e}=20$ Myr and $\tau_{e}=5.1$ Myr.
@@
While the giant planet instability model has reproduced many important features of the currently observed structure within the giant planets and TNOs, the evolution in between an instability and the current TNO configuration remains unknown. This includes the possibility that Uranus and Neptune were launched onto orbits with eccentricities as high as 0.4 and damped at various magnitudes and rates depending on the mass density of the planetesimal disk. With this in mind, we run a last suite of simulations similar to those described in Figure 7 and 8 but with Uranus’ eccentricity initialized at values higher than Neptune’s. One such example is shown in Figure 9. In this simulation, Uranus’ and Neptune’s eccentricity are initialized at 0.1 and 0.065, respectively, representing a scenario just before a damping phase. Both planets are artificially damped with the same eccentricity-damping timescale of 5 Myr, but Neptune is forced to damp to 0.023 and Uranus to 0.04. Prescribing the same damping timescale is motivated by a few of our tests, instability simulations that show that the planetesimal number density near Uranus and Neptune is comparable after a few million years post-instability (not shown). In this simulation, Neptune’s effect on Uranus is significant enough to damp Uranus’ eccentricity to a midpoint below 0.02, which is lower than we prescribed (top right panel). We saw three outcomes in the 40 simulations instances we ran: 7 simulations behaved similarly to Figure 9, 3 simulations showed signs of planet interactions and Uranus’ eccentricity ended up near the final eccentricity we prescribed, and the remaining simulations showed chaotic planet evolution. Both of the simulations shown in Figure 8 and 9 show a quick, jumping feature in the semi-major axis and eccentricity at early times, which can be attributed to Uranus and Neptune crossing a mean motion resonance.
@@
Figure 9: Fifty million year simulation (orange dots) initialized with the same parameters as those in the late stage evolution from Figure 6, except Uranus (top row) and Neptune (bottom row) start at eccentricities 0.1 and 0.065, respectively. Neptune is migrated artificially (maroon line, bottom left panel) to 26.48 au with an exponential timescale of 11.9 million years. Both planets’ eccentricities are artificially damped with the same timescale of 5 Myr (solid maroon line). Uranus’ eccentricity damps further than the amount we prescribe (top right panel) due to the secular effects with Neptune.
5 Discusion and Summary
@@
The modifications to REBOUND and Mercury6.2 we present in this paper are helpful for running suites of N-body simulations that can probe a large parameter space of orbital configurations for the purpose of constraining the evolution of the outer solar system. We demonstrate how to evolve particles’ orbital evolutions with arbitrary functions by applying user-defined velocities along with the more common, user-defined accelerations. We show that with a one-planet simulation, all orbital elements can be evolved, as prescribed and independently from each other. Introducing a second planet to the system, adds new gravitational interactions between the planets, which affects the amount of control one has over the artificial evolution of each orbital element. Secular perturbations between the planets produce oscillations in eccentricity and inclination, so the final simulated average $e$ and/or $i$ may be lower or higher than the prescribed final value. Planets may also cross mutual mean motion resonances, inducing jumps which may dominate the orbital behavior over the artificial evolution. It is therefore, important to carefully choose initial conditions and consider how the secular frequencies and amplitudes of the planets compare to the desired scale of change and timescale of change to successfully evolve $e$ and $i$ accordingly. In our simulations, we do not impose the conservation of angular momentum, which is nonphysical, but our primary goal is to mimic the presence of physical processes absent from the simulation (e.g. a planetesimal disk interacting with planets). By evolving the planets’ orbits according to a physically motivated process and keeping the gravitational interactions between planets, we allow ourselves to have more control in the simulation while maintaining the effect of planet interactions.
We demonstrate appropriate uses of this code, emphasizing the need for physically motivated functional forms for a given orbital evolution. We focus on Uranus’ and Neptune’s artificial orbital evolution in the context of reproducing epochs of smoother migration and eccentricity damping phases in a well-known planetary instability simulation (Tsiganis et al., 2005). We choose an exponential function to evolve these parameters and confirm eccentricity damping timescales with dynamical friction derivations appropriate for the planetesimal disk in their simulation. We are able to reproduce two stages of smoother migration and eccentricity damping. Artificial evolution is also useful for outer solar system studies that probe how different giant planet dynamical histories affect the orbital configuration of TNOs. A variety of proposed migration histories within the outer solar system, modeled through numerical simulations, produce different planet and TNO structure (for reviews see Nesvorný, 2018; Gladman & Volk, 2021). In fact, a similar migration history, with minuscule changes in initial conditions, could also produce vastly distinct planet trajectories as is the case for planetary instability numerical models as discussed in Section 4, where chaos is the underlying driver (e.g. see Volk & Malhotra, 2019, for a discussion on the expense of running such simulations). A realistic portrayal of a planetary system in an N-body simulation would require all solar system planets and billions of massive planetesimals. The chaotic nature of this problem is further accentuated by the additional gravitational interactions between each body, making exploration of a large parameter space of initial conditions computationally infeasible. The ability to artificially evolve planets’ orbits is therefore beneficial for investigating how TNOs in mean motion resonance (MMR) with Neptune evolve as a result of various migration histories.
An interesting application of this numerical approach is the ability to isolate the impact of one planet’s eccentricity damping on another planet, in the absence of a planetesimal disk. We find that artificially damping Neptune’s eccentricity indirectly damps Uranus’ eccentricity due to their secular coupling. The extent to which Neptune’s eccentricity damping affects Uranus depends on the damping timescale and change in eccentricity. At late stages of evolution, coupling between the latter two planets is important to consider since the planetesimal disk will have dissipated significantly.
@@
We thank the anonymous reviewer for their comments and suggestions, which led to the improvement of this paper. We gratefully acknowledge funding from NASA Emerging Worlds grant 80NSSC21K0376. AHR thanks the LSST-DA Data Science Fellowship Program, which is funded by LSST-DA, the Brinson Foundation, and the Moore Foundation; her participation in the program has benefited this work. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 1842400. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. KV acknowledges additional funding from NASA grants 80NSSC19K0785 and 80NSSC23K0886. RP acknowledges additional funding from NASA Solar System Observations grant 80NSSC21K0289. We acknowledge use of the lux supercomputer at UC Santa Cruz, funded by NSF MRI grant AST 1828315.
Appendix A Simulation Parameters
The initial conditions for the simulations shown in this paper are provided in Table 1.
simulation initial parameters and imposed evolution parameters
units: au, Myr, deg
Fig. $m$ $a$ $e$ $i$ $\omega$ $\Omega$ $f$
$a_{0}$ $\Delta a$ $\tau_{a}$ Eq $e_{0}$ $\Delta e$ $\tau_{e}$ Eq $i_{0}$ $\Delta i$ $\tau_{i}$ Eq $\omega_{0}$ $\Delta\omega$ $\tau_{\omega}$ Eq $\Omega_{0}$ $\Delta\Omega$ $\tau_{\Omega}$ Eq $f_{0}$
1 $M_{J}$ 5.2 1.8 10 7a 0.2 -0.1 5 7c 10 5 4 7b 50 35 80 7d 30 60 20 7b 240
2 $M_{J}$ 6 -1 10 7c 0.2 -0.1 5 7c 5 -2 20 7c 50 - - - 30 - - - 240
$M_{N}$ 23 7 10 7c 0.1 0.2 5 7c 10 -8 20 7c 200 - - - 280 - - - 250
3 $M_{J}$ 6 - - - 0.2 - - - 10 5 20 7b 50 - - - 30 - - - 240
a $M_{N}$ 23 - - - 0.1 - - - 20 -15 20 7b 260 - - - 287 - - - 277
b $M_{J}$ 6 - - - 0.2 - - - 10 5 5 7b 50 - - - 30 - - - 240
$M_{N}$ 23 - - - 0.1 - - - 20 -15 5 7b 260 - - - 287 - - - 277
c $M_{J}$ 6 - - - 0.2 - - - 10 2 20 7b 50 - - - 30 - - - 240
$M_{N}$ 23 - - - 0.1 - - - 20 -5 20 7b 260 - - - 287 - - - 277
4 $M_{J}$ 6 - - - 0.05 - - - 1 - - - 50 100 50 7d 30 - - - 240
o $M_{N}$ 23 - - - 0.05 - - - 1 - - - 200 - - - 70 - - - 270
b $M_{J}$ 6 - - - 0.05 - - - 1 - - - 50 - - - 30 - - - 240
$M_{N}$ 23 - - - 0.05 - - - 1 - - - 200 - - - 70 - - - 270
6 $M_{J}$ 5.15 - - - 0.052 - - - 2 - - - 164 - - - 163 - - - 285
$M_{S}$ 8.75 - - - 0.1 - - - 4 - - - 308 - - - 247 - - - 73
$M_{U}$ 18.6 0.7 7.2 7c 0.06 -0.042 1.5 7c 6 - - - 179 - - - 276 - - - 243
$M_{N}$ 23.7 1.7 11 7c 0.06 -0.04 1.5 7c 5 - - - 126 - - - 10 - - - 90
$M_{J}$ 5.15 - - - 0.052 - - - 2 - - - 187 - - - 151 - - - 36
$M_{S}$ 8.75 - - - 0.1 - - - 4 - - - 89 - - - 264 - - - 120
$M_{U}$ 19.3 0.4 17 7c 0.054 -0.024 20 7c 6 - - - 220 - - - 308 - - - 133
$M_{N}$ 25.5 0.93 9.6 7c 0.055 -0.033 6 7c 5 - - - 187 - - - 36 - - - 269
7 $M_{J}$ 5.15 - - - 0.052 - - - 2 - - - 122 - - - 286 - - - 176
$M_{S}$ 8.75 - - - 0.1 - - - 4 - - - 167 - - - 160 - - - 37
$M_{U}$ 19.3 - - - 0.054 - - - 6 - - - 348 - - - 164 - - - 278
$M_{N}$ 25.5 0.93 9.6 7c 0.055 -0.033 6 7c 5 - - - 300 - - - 76 - - - 49
8 $M_{J}$ 5.15 - - - 0.052 - - - 2 - - - 260 - - - 65 - - - 172
$M_{S}$ 8.75 - - - 0.1 - - - 4 - - - 120 - - - 359 - - - 28
$M_{U}$ 19.34 - - - 0.06 - - - 6 - - - 136 - - - 322 - - - 153
$M_{N}$ 25.58 0.9 11.9 7c 0.065 -0.042 5.1 7c 5 - - - 40 - - - 302 - - - 236
9 $M_{J}$ 5.15 - - - 0.052 - - - 1 - - - 73 - - - 260 - - - 161
$M_{S}$ 8.75 - - - 0.1 - - - 2 - - - 284 - - - 319 - - - 286
$M_{U}$ 19.34 - - - 0.1 -0.06 5.1 7c 1 - - - 185 - - - 191 - - - 268
$M_{N}$ 25.58 0.9 11.9 7c 0.065 -0.042 5.1 7c 1 - - - 177 - - - 86 - - - 356
Table 1: Planet parameters—mass ($m$), semimajor axis ($a$), eccentricity ($e$), inclination ($i$), argument of pericenter ($\omega$), longitude of ascending node ($\Omega$), and true anomoly (ff)—for simulations illustrated in each figure. If an imposed evolution was used for a planet’s orbital element, the columns, $\Delta$, $\tau$, and functional form (Eq 7) are shown. A “-” denotes the absence of an imposed evolution. The angles $\omega$, $\Omega$, and $f$ for simulations in Figures 6-9 are produced by a random number generator and are rounded to a whole number here. Because of the chaotic nature of this configuration, i.e. the planets are within the range of another planet’s gravitational influence, several simulations with randomized angles are necessary to have at least one evolution similar to these, as discussed in Section 4. The planet masses $M_{J}$, $M_{S}$, $M_{U}$, $M_{N}$ are Jupiter, Saturn, Uranus, and Neptune masses, respectively. Semimajor axis, angles, and timescales are in units of au, degree, and Myr, respectively.
Appendix B Modifications to the equations of motion